OpenStudy (pokemon23):

Express each product in simplest standard form. (a-4) (a^2-2a+1)

OpenStudy (pokemon23):

omg I asked by 200 question everyone gratz me :)

OpenStudy (unklerhaukus):

gratz

OpenStudy (pokemon23):

explain it to me?

OpenStudy (unklerhaukus):

\[(a-4) (a^2-2a+1)\]\[=(a-4)(a-1)(a-1)\]\[=(a-4)(a-1)^2\]

OpenStudy (anonymous):